# breaklines

### v3

@@ 7.3.2. Translation Bugs
Our artifacts (Lab, 2025) contain partial translations and fixed versions as separate commits. These commits can serve as useful benchmarks for evaluating fault localization, program repair, and test generation techniques. This section shows four instances of such translation bugs.
The two most prevalent sources of translation bugs are mismatches between APIs and behavioral differences in the PLs. The code snippet below demonstrates a bug that happened due to a mismatch in the logic of Calendar (Java) and datetime (Python). Line $3$ in Java sets the MONTH field to 0, which corresponds to the first month of the year (January). Similarly, the Python translation sets the month attribute to 0; however, in the Python library, January is the first month, i.e., the correct translation should use index 1.
@@
The next example shows the difference in implicit type casting between the two Pls. Line $5$ in Java source code concatenates a String with nullStr = null. During execution, Java runtime silently casts null to a String and then performs the binary operation on it. In Python, concatenating an str with None results in a TypeError as the operands of the binary operation has different types. A correct Python translation requires explicit casting of None to str as shown in Line $5$.
@@
The third example shows an instance of write(int b) method from ByteArrayOutputStream class, where the least significant $8$ bits of the integer (b2 << 4) | (b3 >> 2) are directly written to the stream. The incorrect Python translation attempts to construct a bytes object using a singleton list with the input integer before writing it to an object of io.BytesIO. However, this neglects that the bytes() constructor requires the integers in the input iterable to be strictly in the range of [0, 255]. Thereby, a ValueError is thrown when b2 is large. The correct translation requires 0xF, a mask that maintains only the 4 lowest bits of b2 before left-shifting by 4 as shown in Line $3$ under Python translation. Given that b3 and b4 each contain no more than 8 bits, this change ensures the least significant $8$ bits of (b2 << 4) | (b3 >> 2) are correctly written to the BytesIO object.
@@
The last code snippet demonstrates the Java behavior of an iterator that is unavailable in Python. The incorrect Python translation uses next() to implement both .next() and .hasNext() APIs. However, calling next() increments the iterator in Python. The correct translation should implement PeekableIterator interface in Python with a method hasNext() -> bool.
@@
Implications. In order to obtain correct translation, especially for translating APIs, models need to generate test cases as well, which will validate the translated fragment in isolation. This could be an interesting direction for applying the agentic approach, where the orchestrating agent can decide when to generate test cases, and the test case generator agent gets all the information by running static analysis tools, gathering context from previous runs, collecting API documentation by crawling internet, and can finally generating the translation based on all the information.
@@ 7.4. RQ3: Impact of Test Decomposition
We previously showed the effectiveness of test translation in validating the runtime behavior or functional correctness of application method fragments (§7.2.2). To better understand how test decomposition helps test translation coupling effect, we collected translated unit tests with at least two corresponding decomposed test fragments. We further filtered out these unit tests and kept those that all their decomposed test fragments were executed, regardless of passing or failing. The yellow bars in Figure 6 show the percentage of our selected unit tests from the translated test suites. None of the tests met the criteria for Commons-CSV, so we excluded it from further investigation.
@@
We categorized the selected unit tests into two groups: Those with all their test fragments pass (green bars in Figure 6), and those with at least one test fragment fail (red bars in Figure 6). For the unit tests in the latter group, we calculate the pass rate of the decomposed test fragments. The blue box chart in Figure 6 shows the distribution of the measured pass rate per unit test. These unit tests would have been marked as fail without test decomposition. However, we can observe that these tests can be decomposed into test fragments that $62.41\%$ of them pass. These results confirm how decomposed test fragments were useful in helping developers localize translation bugs more easily and resolve the translation bugs faster.
Summary. Test decomposition unburdens validation of application method fragments from incorrect translations. $62.41\%$ of test fragments for unit tests that would have been marked as failed achieved a test pass.
@@ 7.5. RQ4: Impact of Test Coverage
While existing developer-written tests are useful for validating functional equivalence, they can pose two major issues for automated code translation and validation. First, the coverage for these tests can be extremely low (e.g., $13.02\%$ for Commons-FileUpload (Foundation, 2024e) reported in Table 1), preventing most of the code from being validated (as shown in RQ1, the translation validation rate strongly correlates with test suites’ (method) coverage). Second, executing a developer-written test can have a long call sequence. To show the positive impact of better, more focused tests with higher coverage on translation validation, we automatically generated additional tests using EvoSuite (Fraser and Arcuri, 2011). We used DynaMOSA (Panichella et al., 2018) as an optimization algorithm and set a timeout of $120$ seconds as a stopping criterion for the generation of evolutionary tests.
@@
Table 3 compares the developer-written and EvoSuite tests’ properties. Evosuite tests cover more methods than the developer-written test suite ($66.87\%$ Method Coverage compared to $56.57\%$). Furthermore, the average number of methods executed per single test is almost half that of decomposed test suites ($11.41$ compared to $21.84$ methods). To demonstrate the impact of test quality on AlphaTrans’s overall performance, we translated and executed the EvoSuite-generated tests. Corroborated by the numbers under TPR+ and ATP+ columns, we can see that augmenting the test suite increases the method coverage, and thereby, the TPR and ATP numbers from RQ1 by $5.85\%$ and $2.11\%$, respectively. Not all the EvoSuite tests have assertions, and even if they do, the quality of the assertions could be lower compared to developer-written tests (e.g., checking trivial properties that result in a test pass). Thereby, we only claim that higher TPR and ATP of such tests enhance runtime validation, which is still promising in code translation. EvoSuite is incompatible with Java $21$, and hence GraalVM, preventing us from using Evosuite-generated tests in AlphaTrans. We anticipate that including it in the loops could have also improved the overall quality of translations.
@@ 7.6. RQ5: Ablation Study
Subjects AMF Syntax Check (%) SNEF (%) GS (%) GF (%) GE (%) TNEF (%) ATP (%) OTF (%) MTF (%) ATF (%) TPR (%)
cli 276 100 5.80 0 0 94.20 85.51 0 2.17 1.09 5.43 0.46
codec 678 98.82 10.62 0 0 89.38 72.86 2.06 5.46 2.80 6.19 5.04
csv 235 100 8.51 0 0 91.49 88.09 0.43 0.43 0.85 1.70 0.32
exec 253 100 46.25 26.88 4.35 22.53 49.41 1.58 1.19 0 1.58 4.29
fast-pfor 754 98.14 46.02 0 0 53.98 38.20 0.93 6.63 0.80 7.43 4.88
fileupload 168 100 85.12 10.71 0.60 3.57 8.93 1.19 2.98 1.19 0.60 23.08
graph 556 99.82 40.65 0 0 59.35 56.12 0 1.98 0.36 0.90 5.48
jansi 314 99.68 69.75 0 0 30.25 29.94 0 0.32 0 0 0
pool 574 99.83 70.56 0 0 29.44 27.00 1.39 0 0.35 0.70 5.48
validator 623 99.20 33.39 18.78 20.06 27.77 66.29 0 0 0.32 0 0.43
Total 4431 99.32 40.01 4.58 3.09 52.31 52.79 0.81 2.57 0.86 2.96 3.05
@@ 7.6.2. Choice of LLM
For this experiment, we replaced the DeepSeek-Coder-33b-Instruct with GPT-4o and repeated the entire pipeline of AlphaTrans (Table 5). A stronger model such as GPT-4o improves the translation quality —functional equivalence increases from $25.14\%$ to $27.95\%$. For some projects, the ATP and TPR rates are higher for DeepSeek-Coder-33b-Instruct translations, while for the others, GPT-4o results in higher values. We investigated each LLM’s successful AMF translations to better understand the differences. We observed a huge overlap between successful AMFs and the unique benefits each LLM provides in code translation (Figure 7). GPT-4o handles API translation and type casting better, resolving the first three translation bugs discussed in §7.3. In contrast, it tends to add unnecessary code, mostly due to error handling, which results in a functional mismatch. In the example below, create2 method can take a None variable, and its implementation performs error handling when necessary. GPT-4o adds unnecessary error handling code, interfering with program logic and resulting in test failures.
@@
Figure 7. Functional equivalence intersection of AMFs (Graal Success) between LLMs.
@@ 7.6.3. Impact of Program Decomposition
For this ablation study, we prompted GPT-4o and DeepSeek-Coder-33b-Instruct file-by-file and evaluated translation correctness through test execution and GraalVM (Table 6). Not surprisingly, a considerable percentage of the files exceeded the model context window size, particularly for DeepSeekCoder, with $9.08\%$ of the files encountering this problem. Among the prompted files, $21.36\%$ ($19.44\%$ for DeepSeekCoder and $1.92\%$ for GPT-4o) were syntactically incorrect. Translations that passed the syntactic correctness check did not pass any translated test execution, whereas Graal validated $2.56\%$ of the files in total ($1.02\%$ for DeepSeekCoder and $1.54\%$ for GPT-4o). Note that file-level Graal validation does not mean that all the methods are correctly translated and validated; only the methods in the class are executed by tests. Manual analysis of the results of this experiment revealed that one prominent culprit of test failures was LLM hallucination with method/variable names. Given that we had no skeleton construction in this baseline, such issues could not be avoided; this demonstrates the usefulness of skeleton construction and incremental translation in AlphaTrans.
@@ 8. Related Work
There are generally two main approaches to translating code from one programming language to another: using transpilers and statistical machine translation and leveraging language models.
8.1. Code translation using non-LLM-based approaches.
In this domain, tools like C2Rust (Immunant, 2024), CxGo (Transpile, 2024), Sharpen (Project, 2023), and Java2CSharp (Irwin, 2024) have been developed to translate code from C to Rust, C to Go, and Java to C# respectively. However, a recent study (Pan et al., 2024) revealed that, apart from C2Rust and CxGo, other tools lack proper maintenance. For CxGo, language models outperform traditional approaches, whereas for C2Rust, language models generate safer but less effective code, aligning with the primary goal of translating C code to Rust. In terms of statistical machine translation, works by Nguyen  et al. (Nguyen et al., 2013, 2014, 2015), and Chen et al. (Chen et al., 2018) focus on translating Java to C#. Additionally, deep learning approaches have been utilized for code translation (Roziere et al., 2020, 2021). However, none of these efforts have tackled the translation of real-world Java projects to Python.
@@
8.2. Code translation using LLMs.
Recently, large language models have been employed for code translation (Pan et al., 2024; Tipirneni et al., 2024; Di et al., 2024; Yan et al., 2023; Yin et al., 2024; Jiao et al., 2023), demonstrating high success rates on crafted examples but poor performance on real-world projects. Other studies (Zhu et al., 2022; Ahmad et al., 2023) have also utilized language models for code translation, mainly focusing on crafted benchmarks. Recently, there have been works that use transpiler output to guide the code translation (Yang et al., 2024b). However, the limitation of such work is the availability of robust and well-maintained transpilers, which, in many cases, may not be a feasible solution. Nitin et al. (Nitin et al., 2024) introduced a specification-based translation, where natural language specification has been captured from the source code, which helps the translation process. Whereas Yang et al. (Yang et al., 2024a) used tests to assist the translation. However, compared to previous works, the major differences are (a) the first attempt to translate a real-world project, (b) modular translation, and (c) a validation-guided translation approach.
@@ 9. Threats to Validity
Like most approaches, AlphaTrans possesses some limitations and comes with a list of threats to the validity. In this section, we will discuss how we mitigated various threats.
External Validity. One of the key external threats is the generalizability of AlphaTrans. We built and evaluated the first version of AlphaTrans to translate from Java to Python. However, our pipeline is generic, and with minimal effort, the current implementation can translate Java programs to more target languages (e.g., languages supported by GraalVM). Furthermore, the majority of the tools that we used support a large set of programming languages such as JavaScript, Ruby, C/C++, and Rust. Another threat is that AlphaTrans removes certain third-party libraries before translation. Translating repository-level code along with their libraries, especially when the target language does not have equivalent libraries, is a challenging problem that we plan to address in our future work.
Internal Validity. One threat can be the manual validation of the translated types. To address that, several authors have verified the types individually and consulted API documents when necessary. Another threat is that, while all successes reported by the GraalVM validation are true successes, we may have underestimated the capabilities of AlphaTrans by a considerable margin due to the significant proportion of errors caused by limitations in the GraalVM validation approach. To mitigate this threat, we manually augment the universal type map to support a comprehensive translation of types. That said, we have implemented driver code templates to provide a mechanism for adding the support for more types by the users of AlphaTrans if needed.
Construct Validity. In order to mitigate construct validity, AlphaTrans is built and validated with well-vetted tools like, GraalVM (Oracle, 2024), JaCoCo (Team, 2024a), Python coverage (Batchelder, 2024), CodeQL (GitHub, 2024), etc.

### v5

@@ 7.3.2. Translation Bugs
Our artifacts (Lab, 2025a) contain partial translations and fixed versions as separate commits. These commits can serve as useful benchmarks for evaluating fault localization, program repair, and test generation techniques. This section shows four instances of such translation bugs.
The two most prevalent sources of translation bugs are mismatches between APIs and behavioral differences in the PLs. The code snippet below demonstrates a bug that happened due to a mismatch in the logic of Calendar (Java) and datetime (Python). Line $3$ in Java sets the MONTH field to 0, which corresponds to the first month of the year (January). Similarly, the Python translation sets the month attribute to 0; however, in the Python library, January corresponds to value 1 for month.
@@
The next example shows the difference in implicit type casting between the two Pls. Line $5$ in Java source code concatenates a String value with null. During execution, Java runtime silently casts null to a String and then performs the concatenation operation on it. In Python, concatenating an str with None results in a TypeError as the operands of the binary operation has different types. A correct Python translation requires explicit casting of None to str as shown in Line $5$.
@@
The third example shows an instance of write(int b) method from ByteArrayOutputStream class, where the least significant $8$ bits of the integer (b2 << 4) | (b3 >> 2) are directly written to the stream. The incorrect Python translation attempts to construct a bytes object using a singleton list with the input integer before writing it to an object of type io.BytesIO. However, this neglects that the bytes() constructor requires the integers in the input iterable to be strictly in the range of [0, 255]. Thereby, a ValueError is thrown when b2 is large. The correct translation requires 0xF, a mask that maintains only the 4 lowest bits of b2 before left-shifting by 4 as shown in Line $3$ under Python translation. Given that b3 and b4 each contain no more than 8 bits, this change ensures the least significant $8$ bits of (b2 << 4) | (b3 >> 2) are correctly written to the BytesIO object.
@@
The last code snippet demonstrates the Java behavior of an iterator that is unavailable in Python. The incorrect Python translation uses next() to implement both next() and hasNext() methods of the Java java.util.Iterator type. However, calling next() increments the iterator in Python. The correct translation should implement PeekableIterator interface in Python with a method hasNext() -> bool.
@@
Implications. To obtain correct translation, especially for code using library APIs, models need to generate test cases as well that can validate the translated fragment in isolation. This could be an interesting direction for applying an agentic approach, where the orchestrating agent can decide when to generate test cases, and the test case generator agent gets all the information by running static analysis tools, gathering context from previous runs, collecting API documentation by crawling the internet, and finally generating the translation based on all the information.
@@ 7.4. RQ3: Impact of Test Decomposition
We previously showed the effectiveness of test translation in validating the runtime behavior or functional correctness of translated method fragments (§7.2.2). To better understand how test decomposition helps address test translation coupling effect, we collected translated unit tests with at least two decomposed test fragments. We further applied filtering and kept only those tests for which all their decomposed fragments were executed, regardless of passing or failing outcomes. The yellow bars in Figure 6 show the percentage of selected unit tests from the translated test suites. None of the tests met the criteria for Commons-CSV, so we excluded it from further investigation.
@@
We categorized the selected unit tests into two groups: those with all passing test fragments (green bars in Figure 6) and those with at least one failing test fragment (red bars in Figure 6). For the unit tests in the latter group, we calculated the pass rate of the decomposed test fragments. The blue box chart in Figure 6 shows the distribution of the measured pass rate per unit test. These unit tests would have been marked as fail without test decomposition. However, we can observe that these tests can be decomposed into test fragments such that $62.41\%$ of them pass. These results indicate how decomposed test fragments were useful in helping developers localize translation bugs more easily and resolve translation bugs faster.
Summary. Test decomposition unburdens validation of translated method fragments from incorrect translations. $62.41\%$ of test fragments for unit tests that would have been marked as failed achieved passing outcomes.
@@ 7.5. RQ4: Impact of Test Coverage
Although existing developer-written tests are useful for checking functional equivalence, they can pose two major issues for automated code translation and validation. First, the coverage for these tests can be extremely low (e.g., $13.02\%$ for Commons-FileUpload (Foundation, 2024e) as reported in Table 1), preventing most of the code from being validated; as our investigation of RQ1 showed, the translation validation rate strongly correlates with test suites’ (method) coverage. Second, a developer-written test can have a long call sequence. To show the positive impact of more-focused tests with higher coverage on translation validation, we automatically generated additional tests using EvoSuite (Fraser and Arcuri, 2011). We used default tool configuration with a time budget of $120$ seconds per class.
@@
Table 3 compares properties of the developer-written and EvoSuite-generated tests. Evosuite tests cover more methods than the developer-written test suite ($66.87\%$ Method Coverage compared to $56.57\%$). Furthermore, the average number of methods executed per single test is almost half that of decomposed test suites ($11.41$ compared to $21.84$ methods). To demonstrate the impact of test quality on AlphaTrans’s overall performance, we translated and executed the EvoSuite-generated tests. Corroborated by the numbers under TPR+ and ATP+ columns, we can see that augmenting the test suite increases the method coverage, and thereby, the TPR and ATP numbers from RQ1 by $5.85\%$ and $2.11\%$, respectively. Not all EvoSuite tests have assertions, and even if they do, the quality of the assertions could be lower compared to developer-written tests (e.g., checking trivial properties with weak fault detection ability). Nonetheless, the higher TPR and ATP of such tests enhance runtime validation, which is still promising in code translation. EvoSuite is incompatible with Java $21$, and hence GraalVM, which prevented us from using Evosuite-generated tests in AlphaTrans. We anticipate that incorporating it in AlphaTrans could improve the overall quality of translations.
@@ 7.6. RQ5: Ablation Study
Subjects AMF SNEF (%) GS (%) GF (%) GE (%) TNEF (%) ATP (%) OTF (%) MTF (%) ATF (%) TPR (%)
cli 276 5.80 0 0 94.20 85.51 0 2.17 1.09 5.43 0.46
codec 678 10.62 0 0 89.38 72.86 2.06 5.46 2.80 6.19 5.04
csv 235 8.51 0 0 91.49 88.09 0.43 0.43 0.85 1.70 0.32
exec 253 46.25 26.88 4.35 22.53 49.41 1.58 1.19 0 1.58 4.29
fast-pfor 754 46.02 0 0 53.98 38.20 0.93 6.63 0.80 7.43 4.88
fileupload 168 85.12 10.71 0.60 3.57 8.93 1.19 2.98 1.19 0.60 23.08
graph 556 40.65 0 0 59.35 56.12 0 1.98 0.36 0.90 5.48
jansi 314 69.75 0 0 30.25 29.94 0 0.32 0 0 0
pool 574 70.56 0 0 29.44 27.00 1.39 0 0.35 0.70 5.48
validator 623 33.39 18.78 20.06 27.77 66.29 0 0 0.32 0 0.43
Total 4431 40.01 4.58 3.09 52.31 52.79 0.81 2.57 0.86 2.96 3.05
@@ 7.6.2. Choice of LLM
For this experiment, we replaced the DeepSeek-Coder-33b-Instruct with GPT-4o and repeated the entire pipeline of AlphaTrans (Table 5). A stronger model such as GPT-4o improves the translation quality—functional equivalence increases from $25.14\%$ to $27.95\%$. For some projects, the ATP and TPR rates are higher for DeepSeek-Coder-33b-Instruct translations, whereas for the others, GPT-4o results in higher values. We investigated each LLM’s successful AMF translations to better understand the differences. We observed a huge overlap between successful AMFs and the unique benefits each LLM provides in code translation (Figure 7). GPT-4o handles API translation and type casting better, resolving the first three translation bugs discussed in §7.3. In contrast, it tends to add unnecessary code, mostly due to error handling, which results in a functional mismatch. In the example below, create2 method can take a None value, and its implementation performs error handling when necessary. GPT-4o adds unnecessary error handling code, interfering with program logic and resulting in test failures.
@@ 7.6.3. Impact of Program Decomposition
For this ablation study, we prompted GPT-4o and DeepSeek-Coder-33b-Instruct file-by-file and evaluated translation correctness through test execution and GraalVM (Table 6). Not surprisingly, a considerable percentage of the files exceeded the model context window size, particularly for DeepSeekCoder, with $9.08\%$ of the files encountering this problem. Among the prompted files, $21.36\%$ ($19.44\%$ for DeepSeekCoder and $1.92\%$ for GPT-4o) were syntactically incorrect. Translations that passed the syntactic correctness check did not pass any translated test execution, whereas GraalVM validated $2.56\%$ of the files in total ($1.02\%$ for DeepSeekCoder and $1.54\%$ for GPT-4o). Note that file-level GraalVM validation does not mean that all the methods are correctly translated and validated—only the methods in the class that are executed by tests. Manual analysis of the results of this experiment revealed that one prominent culprit of test failures was LLM hallucination with method/variable names. Given that we had no skeleton construction in this baseline, such issues could not be avoided; this demonstrates the usefulness of skeleton construction and incremental translation in AlphaTrans.
@@
Figure 7. Functional equivalence overlap of AMFs (Graal Success) between LLMs.
@@ 8. Related Work
There are generally two main categories of techniques for translating code from one PL to another: (1) using transpilers and statistical machine translation and (2) leveraging language models.
8.1. Code Translation Using Non-LLM-Based Approaches
Tools like C2Rust (Immunant, 2024), CxGo (Transpile, 2024), Sharpen (Project, 2023), and Java2CSharp (Irwin, 2024) translate code from C to Rust, C to Go, and Java to C# respectively. A series of statistical machine translation techniques (Nguyen et al., 2013a, b, 2015; Chen et al., 2018) focus on translating Java to C#. Deep learning approaches have also been applied for code translation (Roziere et al., 2020; Rozière et al., 2021). None of these efforts have tackled translating real-world Java projects to Python. LLM-based techniques are also superior to transpilers in terms of performance or readability (Pan et al., 2024).
@@ 8.1. Code Translation Using Non-LLM-Based Approaches
8.2. Code Translation Using LLMs
Recently, LLMs have been employed for code translation (Pan et al., 2024; Tipirneni et al., 2024; Di et al., 2024; Yan et al., 2023; Yin et al., 2024; Jiao et al., 2023), demonstrating high success rates on crafted examples but poor performance on real-world projects. Other studies (Zhu et al., 2022; Ahmad et al., 2023) have also applied language models for code translation, mainly focusing on crafted benchmarks. Concurrently to our work, two other techniques for repository-level code translation focusing on different language pairs were proposed (Shetty et al., 2024; Zhang et al., 2024). Syzygy (Shetty et al., 2024) translates repository-level C to Rust using GPT-4. Oxidizer (Zhang et al., 2024) leverages language feature mapping and type-driven techniques for translating Go to Rust. Both techniques use I/O equivalence for validating their translations. There are also approaches that use transpiler output to guide LLM-based code translation (Yang et al., 2024b). However, the limitation of such work is the availability of robust and well-maintained transpilers, which, in many cases, may not be feasible. Nitin et al. (Nitin et al., 2024) presented a specification-based translation, where a natural language specification is captured from the source code, which helps the translation process. Yang et al. (Yang et al., 2024a) used tests to assist the translation. Compared to previous work, our contributions include a compositional and validation-guided code translation approach that leverages two types of validation techniques and evaluation of the approach on 10 real-world projects.
@@ 9. Threats to Validity
External Validity. One of the key external threats is the generalizability of AlphaTrans. We built and evaluated the first version of AlphaTrans to translate from Java to Python. However, our pipeline is generic, and with minimal effort, the current implementation can translate Java programs to more target languages (e.g., languages supported by GraalVM). Furthermore, the majority of the tools that we used support a large set of programming languages such as JavaScript, Ruby, C/C++, and Rust.
Internal Validity. One threat can be the manual validation of the translated types. To address that, several authors verified the types individually and consulted API documents when necessary. Another threat is that, while all successes reported by the GraalVM validation are true successes, we may have underestimated the capabilities of AlphaTrans by a considerable margin due to the significant proportion of errors caused by limitations in the GraalVM validation approach. To mitigate this threat, we manually augmented the universal type map to support a more comprehensive translation of types. That said, we have implemented driver code templates to provide a mechanism for adding the support for more types by the users of AlphaTrans if needed.
Construct Validity. In order to mitigate construct validity, AlphaTrans is built and validated with well-vetted tools, such as GraalVM (Oracle, 2024), JaCoCo (Team, 2024a), Python coverage (Batchelder, 2024), CodeQL (GitHub, 2024), etc.